Assume that a procedure yields a binomial distribution with n=2 trials and a probability of success of p=0.40. Use a binomial probability table to find the probability that the number of successes x is exactly 1
Answer #1

Use the table with n=2 and p = 0.4

q = 1 - p = 0.6

The probability is given by:

^{2}C_{2}p^{2}q^{0}+ ^{2}C_{1} p^{1}q^{1}+^{2}C_{0}p^{0}q^{2}

Any number raised to the power 0 is 1.

This simplifies to:

p^{2}+2*p*q+q^{2} = (04)^{2}+2(0.4)(0.6)+(0.6)^{2}=0.16+0.48+0.36 = 1